how old is a bone if it presently contains 0.3125g of C-14, but it was estimated to have originally contained 80.0g of C-14 (C-14 Half-life=5730 years)
1
Expert's answer
2017-04-24T09:06:07-0400
Amount remaining: 0.3125 g / 80.000 g = 0.00390625 g

(1/2) n = 0.00390625

n log 0.5 = log 0.00390625

The number of half-lives elapsed n = 8

5730 · 8 = 45840 years